# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v1

@@ 4.1 Replication Encoding
    The slots indexed by $S^{c}$ are filled with a same message $m$, and the complement slots indexed by $S$ are filled with $n/2$ random values $v_{i}\overset{\$}{\leftarrow}\mathbb{Z}_{t}$.
We call the slots indexed by $S$ as verification slots, and the slots indexed by $S^{c}$ as computation slots. The server evaluates $f$ independently to the slots. After evaluating $f$ on ciphertexts, its decryption contains $f(v_{i})$’s in the verification slots and $f(m)$ in the computation slots. For verification, it is assumed that the client knows $f(v_{i})$’s and their integrity in advance, either by outsourcing the computation of $f(v_{i})$’s without encryption to a third party using existing verifiable computation techniques or by securely computing them himself. Note that the verification values are not published, unlike the VOPRF scheme in [ADDG24]. The client verifies as follows:
@@
4.2 Attack on REP
Learning or guessing $S\subset\{1,\dots,n\}$ is not easy. Instead of attaining $S$, we take a similar approach to the attack on vADDG: To learn $S$ in encrypted state by computing $\textsf{Enc}(J_{S})$ for $J_{S}:=(j_{1},\dots,j_{n})$ where $j_{i}=0$ if $i\in S$ and $j_{i}=1$ otherwise. Once the adversary recovers $\textsf{Enc}(J_{S})$, the ciphertext can be easily forged by computing the following:
${}_{\textsf{forged}}:=\mathsf{Mult}(\textsf{Enc}(J_{S}),\textsf{Eval}_{g}(\textsf{ct}{}))+\mathsf{Mult}(\textsf{Enc}(J_{S^{c}}),\textsf{Eval}_{f}(\textsf{ct}{})),$ ct ) ) ,
where $f$ is the circuit that the client requested to evaluate, and $g$ is any modified circuit by the adversary. Then $\textsf{ct}{}_{\textsf{forged}}$ provides the client with a malicious result $g(m)$ but successfully passes the verification step.
To recover $\textsf{Enc}(J_{S})$, we designed a circuit to extract the position vector of the common values. Figure 1 shows the functionality of the desired algorithm; if the $i$th slot is chosen, the output of this algorithm is the position vector of the slots equal to this value. We do not consider the case where $v_{i}=m$ for some $i\in S$, as in such case the client would already have attained $f(v_{i})=f(m)$.
@@ 4.2 Attack on REP
Consider $e(x)=1-x^{\phi(t)}$ where $\phi$ is the Euler’s totient function so that $e(x-y)=0$ if $x-y\in\mathbb{Z}_{t}^{\times}$, and otherwise $e(x-y)=1$. If $t=p$, $e(x-y)$ is the desired $\mathsf{Compare}(x,y)$ function. Now suppose that $t=p^{r}$ and define a polynomial $\alpha$ as follows:
$\alpha(x):=e(x)\cdot\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u).$
@@
    If $x$ is a nonzero and non-unit, then $x+1\in\mathbb{Z}_{t}^{\times}\setminus\{1\}$. Thus
    $\alpha(x)$ $=e(x)\left(\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u)\right)$
    $=e(x)\cdot 0=0.$
@@
1:procedure $\mathsf{CVS}$($\textsf{ct}{},i$)
2:     ${}^{\prime}\leftarrow\mathsf{\textsf{Eval}_{Duplicate}}_{i}(\textsf{ct}{})$ ct ) $\triangleright$ $\textsf{ct}{}^{\prime}=\textsf{Enc}((x_{i},\dots,x_{i}))$
3:     ${}_{\textsf{CVS},i}\leftarrow\textsf{Eval}_{\mathsf{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ ct )
4:     return $\textsf{ct}{}_{\mathsf{CVS},i}$
5:end procedure
Lemma 2.
For $i\overset{\$}{\leftarrow}\{1,\dots,n\}$, $\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ in the algorithm 2 outputs $\mathsf{Enc}(J_{S})$ with probability of $|S|/n=1/2$ in $O(\log t)$ homomorphic multiplications and rotations.
@@
Let ct be an encryption of a vector that has $m$ in computational slots $S^{c}$. If $i\in S^{c}$, then ${}^{\prime}\leftarrow\textsf{Eval}_{\mathsf{Duplicate}_{i}}(\textsf{ct}{})$ ct ) is an encryption of $(m,\dots,m)$. Consequently, $\mathsf{Eval_{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ returns $\textsf{Enc}(J_{S})$.
@@
Now we show how to increase the success probability of the attack by repeating Algorithm 2 for different $i$’s. For any fixed $i$, if we sum up the values in the $n$ slots of $\textsf{ct}{}_{\textsf{cvs},i}$, it outputs $|S|$ or $1$ depending on whether $i\in S$ or not. i.e. the server can homomorphically distinguish whether $i\in S$ or not. Using this principle, by repeating the algorithm 2, the adversary can construct a cheating circuit that deterministically outputs $\textsf{Enc}(J_{S})$. See Algorithm 3.
@@
1:procedure $\mathsf{is\_in\_set}_{S}$($\textsf{ct}{},i$)
2:     $\textsf{ct}{}_{\textsf{cvs},i}\leftarrow\textsf{CVS}(\textsf{ct}{},i)$
3:     $\textsf{ct}{}_{\textsf{rs}}\leftarrow\textsf{RotSum}(\textsf{ct}{}_{\textsf{cvs},i})$triangleright $\textsf{ct}{}_{\textsf{rs}}=\textsf{Enc}(1,\dots,1)$ or $\textsf{Enc}(n/2,\dots,n/2)$
4:     $\textsf{ct}{}_{\textsf{bool},i}\leftarrow\textsf{Eval}_{\mathsf{Interpolate}}(\textsf{ct}{}_{\textsf{rs}})$
5:     return $\textsf{ct}{}_{\textsf{bool},i}$triangleright $\textsf{ct}{}_{\textsf{bool},i}=\textsf{Enc}(0,\dots,0)$ or $\textsf{Enc}(1,\dots,1)$
6:end procedure
7:
8:procedure $\mathsf{Recover}_{S}$(ct)
9:     $\textsf{ct}{}_{\textsf{cvs},i}\leftarrow\textsf{CVS}(\textsf{ct}{},i)$
10:     $\textsf{ct}{}_{\textsf{bool},i}\leftarrow\textsf{is\_in\_set}_{S}(\textsf{ct}{},i)$
11:     $\textsf{ct}{}_{S}^{\prime}\leftarrow\sum_{i=1}^{n/2+1}\sf{Mult}(\textsf{ct}{}_{\textsf{bool},i},\textsf{ct}{}_{\textsf{cvs},i})$
12:     $\textsf{ct}{}_{S}\leftarrow\textsf{Eval}_{\textsf{Normalize}_{r}}(\textsf{ct}{}_{S}^{\prime})$
13:     return $\textsf{ct}{}_{S}$
14:end procedure
@@
If the rotation key with index 1 is given, $\mathsf{Recover}_{S}(\mathsf{ct})$ in Algorithm 3 can be evaluated within a computational cost of $O(n\log t)$, deterministically outputting $\textsf{Enc}(J_{S})$.
@@
Since the slots in $\textsf{ct}{}_{\textsf{cvs},i}$ contain $n/2$ ones if $i\in S^{c}$, and a single one if $i\in S$, applying $\textsf{Eval}_{\textsf{RotSum}}$ yields either $\textsf{Enc}\left(\frac{n}{2},\dots,\frac{n}{2}\right)$ or $\textsf{Enc}(1,\dots,1).$ Consequently, by interpolation, the output $\textsf{ct}{}_{\text{bool},i}$ is a vector of Boolean values, either $(1,\dots,1)$ or $(0,\dots,0)$. Thus, for $i=1,\dots,n/2+1$, the sum of the $\textsf{ct}{}_{\textsf{cvs},i}$ after homomorphic multiplication with $\textsf{ct}{}_{\text{bool},i}$ decrypts to the same value as $|S^{c}\cap\{1,\dots,n/2+1\}|\cdot\textsf{Enc}(J_{S}).$ Since $S\subset\{1,\dots,n\}$ is a random subset of size $n/2$, $|S^{c}\cap\{1,\dots,n/2+1\}|$ is nonzero. Thus, after evaluating $\textsf{Normalize}_{1+n/2}$, we obtain $\textsf{Enc}(J_{S})$ with these probabilities.
Furthermore, the procedure $\mathsf{Recover}_{S}(\textsf{ct}{})$ calls $\textsf{CVS}_{i}$ for $i=1,\dots,n/2+1$, which incurs a cost of $O(n\log t)$ homomorphic operations. The cost of the remaining operations does not exceed $O(n\log t)$. ∎
@@
REP is not secure against the presence of $\mathcal{A}_{c}^{\textsf{rtk}{}_{1}}$.
@@
If BFV bootstrapping needs $\textsf{rtk}{}_{1}$, REP is not bootstrappable or insecure against $\mathcal{A}_{c}$.
@@ 4.3 Attack on Replication Encoding with Multiple Secret Keys
We introduce $\textsf{REP}{}^{\sf MSK}$, a countermeasure to the previous deterministic attack. The previous attack evaluates homomorphic rotations to compare values in different slots and recover $\textsf{Enc}(J_{S})$. Notably, a rotation key with index 1 is essential for executing this attack. Meanwhile, in REP, a single value $m$ is encoded into $n$ slots. Therefore, to perform operations on two values $m_{1}$ and $m_{2}$ encoded via REP, only rotation keys whose index is a multiple of $n$ will be required. Nevertheless, we note that a rotation key with index 1 remains essential for the bootstrapping of homomorphic encryption. Thus, the idea of preventing the above attack by not providing a rotation key with index 1 is not adequate.
@@
1:procedure $\mathsf{Extended\_Attack}$($\textsf{ct}{},A$)
2:     ${}_{\textsf{bool}}\leftarrow\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$ ct )
3:     $\mathsf{pt}_{\textsf{Masking}}\leftarrow(\overbrace{1,\dots,1}^{n},\overbrace{0,\dots,0}^{N-n})$
4:     $\textsf{ct}{}_{S}\leftarrow\textsf{CMult}(\textsf{ct}{}_{\sf bool},\mathsf{pt}_{\sf Masking})$
5:     return $\textsf{ct}{}_{S}$
6:end procedure
@@
Let $A$ be a randomly chosen subset of $\mathbb{Z}_{t}$ with size $|A|=p\cdot t$. For $\textsf{ct}{}_{S}\leftarrow\mathsf{Extended\_Attack}(\textsf{ct}{},A)$ in the Algorithm 4, the probability
$\Pr\left[\mathsf{Dec}(\textsf{ct}{}_{S})=(J_{S^{c}}\|\overbrace{0\dots 0}^{N-n})\right]=:Q(p)$
@@
Nothing happens in the case (1), and the adversary is caught in the case (2). The case (3) is the case of a successful attack. Now represent the probability of case (3) $Q(p)$ in terms of $p$.
$Q(p)=$ $q^{\mu-1}-q^{\mu}$
$=$ $(1-p)^{\mu-1}-(1-p)^{\mu}.$
By simple calculus, one can verify that $\frac{dQ}{dp}(\mu^{-1})=0$ and $Q(p)$ is maximized when $p=\mu^{-1}$:
$Q(\mu^{-1})=(1-\mu^{-1})^{\mu-1}-(1-\mu^{-1})^{\mu}=(1-p)^{(1/p)-1}-(1-p)^{(1/p)}.$
@@
The attack cost in Theorem 4.2 is determined by the evaluation cost of the characteristic function. In the attack on VOPRF, the evaluation was straightforward since extracting encrypted bits from the TFHE ciphertext string was easy. However, in BFV, homomorphically evaluating $\chi_{A}$ for any randomly given subset $A$ with the full support $\mathbb{Z}_{t}$ generally requires $O(\sqrt{t})$ homomorphic operations, which results in an exponential cost with respect to $\log t$. Therefore, it is necessary to devise an efficient method for evaluating a pseudorandom characteristic function over $\mathbb{Z}_{t}$.
@@
To construct an efficient pseudo-random characteristic function, we assume the pseudo-randomness of the distribution of the roots of unity in $\mathbb{Z}_{t}$: For an integer $t$ and $d\mid\phi(t)$, let $U_{t,d}:=\{x\in\mathbb{Z}_{t}^{\times}:x^{\phi(t)/d}=1\}\subset\mathbb{Z}_{t}^{\times}$. We assume that this subset $U_{t,d}$ of size $\phi(t)/d$ is an unbiased sample from a uniform distribution of $(\phi(t)/d)$-combinations from $\mathbb{Z}_{t}^{\times}$. Specifically, for a large odd prime $p$, $U_{p,2}$ is the set of quadratic residues modulo $p$, whose pseudo-randomness is utilized to construct the Legendre PRF [Dam90]. Although the specific assumptions and parameters in here are different from those of Legendre PRF, we again note that it is sufficient for it to be a function that heuristically produces an unbiased uniform sampling.
@@
    First, suppose that $x$ is a non-unit. Then the left-hand side $(\sf LHS)$ is 00. Also as $x$ is a multiple of $p$, the right-hand side $(\sf RHS)$ is also 00 since it would be multiple of $p^{\phi(t)}\equiv 0\pmod{p^{r}}$.
@@
Note that if $t$ is a prime, then the $x^{\phi(t)}$ term is not needed. Also, note that this function can be homomorphically evaluated using $O(\log t)$ homomorphic multiplications with a multiplication depth of $\lceil\log\phi(t)\rceil$.
@@
$\sf REP^{\sf MSK}$ is not secure against the presence of $\mathcal{A}_{m}.$

### v2

@@ 4.1 Replication Encoding
    The slots indexed by $S^{c}=\{1,\dots,n\}\setminus S$ are filled with a same message $m$, and the complement slots indexed by $S$ are filled with $n/2$ random values $v_{i}\overset{\$}{\leftarrow}\mathbb{Z}_{t}$.
We call the slots indexed by $S$ as verification slots, and the slots indexed by $S^{c}$ as computation slots. After evaluating $f$ to the ciphertext, its decryption contains $f(v_{i})$’s in the verification slots and $f(m)$ in the computation slots. For verification, it is assumed that the client knows $f(v_{i})$’s and their integrity in advance, either by outsourcing the computation of $f(v_{i})$’s without encryption to a third party using existing verifiable computation techniques or by securely computing them himself. Note that the verification values are not published, unlike the VOPRF scheme in [ADDG24]. The client verifies as follows:
@@
4.2 Attacks on REP
Learning or guessing $S\subset\{1,\dots,n\}$ is not easy. Instead of recovering $S$, we take a similar approach to the attack on vADDG: To learn $S$ in encrypted state by computing $\textsf{Enc}(J_{S})$ for $J_{S}:=(j_{1},\dots,j_{n})$ where $j_{i}=1$ if $i\in S$ and $j_{i}=0$ otherwise. Once the adversary recovers $\textsf{Enc}(J_{S})$, the ciphertext can be easily forged by computing the following:
${}_{\textsf{forged}}:=\mathsf{Mult}(\textsf{Enc}(J_{S}),\textsf{Eval}_{f}(\textsf{ct}{}))+\mathsf{Mult}(\textsf{Enc}(J_{S^{c}}),\textsf{Eval}_{g}(\textsf{ct}{})),$ ct ) ) , (1)
where $f$ is the circuit that the client requested to evaluate, and $g$ is any modified circuit by the adversary. (Note that $\textsf{Enc}(J_{S^{c}})=\mathbf{1}^{N}-\textsf{Enc}(J_{S})$, thus computing $\textsf{Enc}(J_{S^{c}})$ is equivalent to computing $\textsf{Enc}(J_{S})$.)
Then $\textsf{ct}{}_{\textsf{forged}}$ provides the client with a malicious result $g(m)$ but successfully passes the verification step.
To recover $\textsf{Enc}(J_{S^{c}})$, we designed a circuit to extract the position vector of the common values. Figure 2 shows the functionality of the desired algorithm; If the $i$th slot is chosen, the output of this algorithm is the position vector of the slots equal to this value. We do not consider the case where $v_{i}=m$ for some $i\in S$, as in such case the client would already have attained $f(v_{i})=f(m)$.
@@ 4.2 Attacks on REP
Consider $e(x)=1-x^{\phi(t)}$ where $\phi$ is the Euler’s totient function so that $e(x-y)=0$ if $x-y\in\mathbb{Z}_{t}^{\times}$, and otherwise $e(x-y)=1$. If $t=p$, $e(x-y)$ is the desired $\mathsf{Compare}(x,y)$ function. Now suppose that $t=p^{r}$ and define a polynomial $\alpha$ as $\alpha(x):=e(x)\cdot\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u).$
@@
    If $x$ is a nonzero and non-unit, then $x+1\in\mathbb{Z}_{t}^{\times}\setminus\{1\}$. Thus $\alpha(x)=e(x)\left(\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u)\right)=e(x)\cdot 0=0.$
@@
1:Input: A ciphertext ct encrypted via REP and an index $i$.
2:Output: A ciphertext $\textsf{ct}{}_{\mathsf{CVS},i}$, possibly an encryption of the vector $J_{S^{c}}$ if $i\in S^{c}$.
3:procedure $\mathsf{CVS}$($\textsf{ct}{},i$)
4:     ${}^{\prime}\leftarrow\mathsf{\textsf{Eval}_{Duplicate}}_{i}(\textsf{ct}{})$ ct ) $\triangleright$ $\textsf{ct}{}^{\prime}=\textsf{Enc}((x_{i},\dots,x_{i}))$
5:     ${}_{\textsf{CVS},i}\leftarrow\textsf{Eval}_{\mathsf{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ ct )
6:     return $\textsf{ct}{}_{\mathsf{CVS},i}$
7:end procedure
Theorem 4.1.
For $i\overset{\$}{\leftarrow}\{1,\dots,n\}$, $\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ in the algorithm 2 outputs $\mathsf{Enc}(J_{S^{c}})$ with probability of $|S|/n=1/2$ in $O(\log t)$ homomorphic multiplications and rotations.
@@
Let ct be an encryption of a vector that has $m$ in computational slots $S^{c}$. If $i\in S^{c}$, then ${}^{\prime}\leftarrow\textsf{Eval}_{\mathsf{Duplicate}_{i}}(\textsf{ct}{})$ ct ) is an encryption of $(m,\dots,m)$. Consequently, $\mathsf{Eval_{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ returns $\textsf{Enc}(J_{S^{c}})$.
@@
Thus, REP is not sound in the presence of the malicious adversary with rotation key with index 1.
Corollary 2.
REP is not sound in the presence of $\mathcal{A}_{\sf mal}^{\sf rtk_{1}}$.
If the attack fails, (i.e. $i\in S$) then the output of $\textsf{ct}{}_{\sf forged}$ of 1 where $\textsf{Enc}(J_{S^{c}})\leftarrow\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ is detected by the client. Thus this attack does not show that the REP is not sound in the presence of the covert adversary.
4.2.1 Increasing Attack Success Probability.
Now we show how to increase the success probability of the attack by repeating Algorithm 2 for different $i$’s. For any fixed $i$, if we sum up the values in the $n$ slots of $\textsf{ct}{}_{\textsf{cvs},i}$, it outputs $|S|$ or $1$ depending on whether $i\in S$ or not. i.e. the server can homomorphically distinguish whether $i\in S$ or not. Using this principle, by repeating the algorithm 2, the adversary can construct a cheating circuit that deterministically outputs $\textsf{Enc}(J_{S^{c}})$. See Algorithm 3.
@@
1:Input: A ciphertext ct encrypted via REP and an index $i$.
2:Output: A ciphertext $\textsf{ct}{}_{\textsf{bool},i}$ which is an encryption of $0^{N}$ if $i\in S$ and $1^{N}$ otherwise.
3:procedure $\mathsf{is\_in\_set}_{S}$($\textsf{ct}{},i$)
4:     $\textsf{ct}{}_{\textsf{cvs},i}\leftarrow\textsf{CVS}(\textsf{ct}{},i)$
5:     $\textsf{ct}{}_{\textsf{rs}}\leftarrow\textsf{RotSum}(\textsf{ct}{}_{\textsf{cvs},i})$triangleright $\textsf{ct}{}_{\textsf{rs}}=\textsf{Enc}(1,\dots,1)$ or $\textsf{Enc}(n/2,\dots,n/2)$
6:     $\textsf{ct}{}_{\textsf{bool},i}\leftarrow\textsf{Eval}_{\mathsf{Interpolate}}(\textsf{ct}{}_{\textsf{rs}})$
7:     return $\textsf{ct}{}_{\textsf{bool},i}$triangleright $\textsf{ct}{}_{\textsf{bool},i}=\textsf{Enc}(0,\dots,0)$ or $\textsf{Enc}(1,\dots,1)$
8:end procedure
9:
10:Input: A ciphertext ct encrypted via REP and an index $i$, an integer $k$.
11:Output: A ciphertext ct possibly encrypting $J_{S^{c}}$
12:procedure $\mathsf{Recover}_{S^{c}}$($\textsf{ct}{},k$) $\triangleright$ Deterministically outputs $\textsf{Enc}(J_{S^{c}})$ when $k=n/2+1$.
13:     $\textsf{ct}{}_{\textsf{cvs},i}\leftarrow\textsf{CVS}(\textsf{ct}{},i)$
14:     $\textsf{ct}{}_{\textsf{bool},i}\leftarrow\textsf{is\_in\_set}_{S}(\textsf{ct}{},i)$
15:     $\textsf{ct}{}^{\prime}\leftarrow\sum_{i=1}^{k}\sf{Mult}(\textsf{ct}{}_{\textsf{bool},i},\textsf{ct}{}_{\textsf{cvs},i})$
16:     $\textsf{ct}{}_{S^{c}}\leftarrow\textsf{Eval}_{\textsf{Normalize}_{k}}(\textsf{ct}{}^{\prime})$
17:     return $\textsf{ct}{}_{S^{c}}$
18:end procedure
@@
If the rotation key with index 1 is given, $\mathsf{Recover}_{S^{c}}(\mathsf{ct},k)$ in Algorithm 3 can be evaluated within a computational cost of $O(k\log t)$, outputting $\textsf{Enc}(J_{S^{c}})$ with probability
$1-\frac{{n/2\choose k}}{{n\choose k}}.$
Otherwise, it outputs an encryption of zero vector $\textsf{Enc}(0,\dots,0)$.
@@
Since the slots in $\textsf{ct}{}_{\textsf{cvs},i}$ contain $n/2$ ones if $i\in S^{c}$, and a single one if $i\in S$, applying $\textsf{Eval}_{\textsf{RotSum}}$ yields either $\textsf{Enc}\left(\frac{n}{2},\dots,\frac{n}{2}\right)$ or $\textsf{Enc}(1,\dots,1).$ Consequently, by interpolation, the output $\textsf{ct}{}_{\text{bool},i}$ is a vector of Boolean values, either $(1,\dots,1)$ or $(0,\dots,0)$. Thus, for $i=1,\dots,k$, the sum of the $\textsf{ct}{}_{\textsf{cvs},i}$ after homomorphic multiplication with $\textsf{ct}{}_{\text{bool},i}$ decrypts to the same value as $|S^{c}\cap\{1,\dots,k\}|\cdot\textsf{Enc}(J_{S^{c}}).$ Since $S\subset\{1,\dots,n\}$ is a random subset of size $n/2$, $|S^{c}\cap\{1,\dots,k\}|$ is zero with probability ($\frac{{n/2\choose k}}{{n\choose k}}$ and otherwise nonzero. Thus, after evaluating $\textsf{Normalize}_{k}$, we obtain $\textsf{Enc}(0,\dots,0)$ with probability ($\frac{{n/2\choose k}}{{n\choose k}}$, or $\textsf{Enc}(J_{S^{c}})$ with probability $1-\frac{{n/2\choose k}}{{n\choose k}}$.
Furthermore, the procedure $\mathsf{Recover}_{S^{c}}(\textsf{ct}{})$ calls $\textsf{CVS}_{i}$ for $i=1,\dots,k$, which incurs a cost of $O(k\log t)$ homomorphic operations. The cost of the remaining operations does not exceed $O(k\log t)$. ∎
By taking $k=n/2+1$, this attack can achieve a $100\%$ success rate [^2] With only negligible probability of failure, due to the properties of the underlying RLWE scheme; see Definitions 2 and 4. Even when $k<n$, the failure probability of the attack is bounded by ($\frac{{n/2\choose k}}{{n\choose k}}\leq 2^{-k}$, which decreases exponentially with $k$. Therefore, if the adversary is willing to trade off success probability for reduced circuit latency, they may choose a sufficiently small value of $k$. Moreover, since the algorithm ${\sf Recover}_{S^{c}}(\textsf{ct}{},k)$ is parallelizable, increasing $k$ may not lead to increased latency, depending on the adversary’s computational capabilities.
Even if the attack fails, the output remains $\textsf{Enc}(0,\dots,0)$, and thus, by the definition of $\textsf{ct}{}_{\sf forged}$ in 1, the attack remains undetected. Hence, the attack can still be carried out by a covert adversary, even when $k$ is sufficiently small.
@@
The scheme $\mathsf{REP}$ is not secure against the covert adversary $\mathcal{A}_{\mathsf{cov}}^{\textsf{rtk}{}_{1}}$.
The following result follows directly from Corollaries 2 and 3.
@@
If BFV bootstrapping requires the key $\textsf{rtk}{}_{1}$, then $\mathsf{REP}$ is either not bootstrappable or insecure against covert/malicious adversaries.
@@ 4.3 Attack on Replication Encoding with Multiple Secret Keys
We introduce $\textsf{REP}{}^{\sf msk}$, a countermeasure to the previous deterministic attack. The previous attack evaluates homomorphic rotations to compare values in different slots and recover $\textsf{Enc}(J_{S^{c}})$. Notably, a rotation key with index 1 is essential for executing this attack. Meanwhile, in REP, a single value $m$ is encoded into $n$ slots. Therefore, to perform operations on two values $m_{1}$ and $m_{2}$ encoded via REP, only rotation keys whose index is a multiple of $n$ will be required. Nevertheless, we note that a rotation key with index 1 remains essential for the bootstrapping of homomorphic encryption. Thus, the idea of preventing the above attack by not providing a rotation key with index 1 is not adequate.
@@
1:Input: A ciphertext ct encrypted via REP and a subset $A\subset\mathbb{Z}_{t}$.
2:Output: A ciphertext possibly encrypting $J_{S^{c}}$ in the first $n$ slots.
3:procedure $\mathsf{Extended\_Attack}$($\textsf{ct}{},A$)
4:     ${}_{\textsf{bool}}\leftarrow\textsf{Eval}_{\chi_{A}}(\textsf{ct}{})$ ct )
5:     $\mathsf{pt}_{\textsf{Masking}}\leftarrow(\overbrace{1,\dots,1}^{n},\overbrace{0,\dots,0}^{N-n})$
6:     $\textsf{ct}{}_{S^{c}}\leftarrow\textsf{CMult}(\textsf{ct}{}_{\sf bool},\mathsf{pt}_{\sf Masking})$
7:     return $\textsf{ct}{}_{S^{c}}$
8:end procedure
@@
Let $A$ be a randomly chosen subset of $\mathbb{Z}_{t}$ with size $|A|=p\cdot t$. For $\textsf{ct}{}_{S^{c}}\leftarrow\mathsf{Extended\_Attack}(\textsf{ct}{},A)$ in the Algorithm 4, the probability
$\Pr\left[\mathsf{Dec}(\textsf{ct}{}_{S^{c}})=(J_{S^{c}}\|\overbrace{0\dots 0}^{N-n})\right]=:Q(p)$
@@
Nothing happens in the case (1), and the adversary is caught in the case (2). The case (3) is the case of a successful attack. Now represent the probability of case (3) $Q(p)$ in terms of $p$: $Q(p)=q^{\mu-1}-q^{\mu}=(1-p)^{\mu-1}-(1-p)^{\mu}.$
By simple calculus, one can verify that $\frac{dQ}{dp}(\mu^{-1})=0$ and $Q(p)$ is maximized when $p=\mu^{-1}$: $Q(\mu^{-1})=(1-\mu^{-1})^{\mu-1}-(1-\mu^{-1})^{\mu}=(1-p)^{(1/p)-1}-(1-p)^{(1/p)}.$
@@
The attack cost in Theorem 4.3 is determined by the evaluation cost of the characteristic function. In the attack on vADDG, the evaluation was straightforward since extracting encrypted bits from the TFHE ciphertext string was easy. However, in BFV, homomorphically evaluating $\chi_{A}$ for any randomly given subset $A$ with the full support $\mathbb{Z}_{t}$ generally requires $O(\sqrt{t})$ homomorphic operations, which results in an exponential cost with respect to $\log t$. Therefore, it is necessary to devise an efficient method for evaluating a pseudorandom characteristic function over $\mathbb{Z}_{t}$.
@@
To construct an efficient pseudo-random characteristic function, we assume the pseudo-randomness of the distribution of the roots of unity in $\mathbb{Z}_{t}$: For an integer $t$ and $d\mid\phi(t)$, let $U_{t,d}:=\{x\in\mathbb{Z}_{t}^{\times}:x^{\phi(t)/d}=1\}\subset\mathbb{Z}_{t}^{\times}$. We assume that this subset $U_{t,d}$ of size $\phi(t)/d$ is an unbiased sample from a uniform distribution of $(\phi(t)/d)$-combinations from $\mathbb{Z}_{t}^{\times}$. Specifically, for a large odd prime $p$, $U_{p,2}$ is the set of quadratic residues modulo $p$, whose pseudo-randomness is utilized to construct the Legendre PRF [Dam90].
@@
    First, suppose that $x$ is a non-unit. Then the left-hand side $(\sf LHS)$ is zero. Also as $x$ is a multiple of $p$, the right-hand side $(\sf RHS)$ is also 00 since it would be multiple of $p^{\phi(t)}\equiv 0\pmod{p^{r}}$.
@@
Note that if $t$ is a prime, then the $x^{\phi(t)}$ term in RHS is not needed. Also, note that this function can be homomorphically evaluated using $O(\log t)$ homomorphic multiplications with a multiplication depth of $\lceil\log\phi(t)\rceil$.
@@
$\sf REP^{\sf msk}$ is not secure against the presence of $\mathcal{A}_{\sf mal}.$